# Exploring Causes of Representational Similarity in Machine Learning Models

### v1

@@ Abstract
Numerous works have noted significant similarities in how machine learning models represent the world, even across modalities. Although much effort has been devoted to uncovering properties and metrics on which these models align, surprisingly little work has explored causes of this similarity. To advance this line of inquiry, this work explores how two possible causal factors—dataset overlap and task overlap—influence downstream model similarity. The exploration of dataset overlap is motivated by the reality that large-scale generative AI models are often trained on overlapping datasets of scraped internet data, while the exploration of task overlap seeks to substantiate claims from a recent work, the Platonic Representation Hypothesis, that task similarity may drive model similarity. We evaluate the effects of both factors through a broad set of experiments. We find that both positively correlate with higher representational similarity and that combining them provides the strongest effect. Our code and dataset are published. 22footnotetext: https://github.com/lizeyu090312/ReprSimCauses 33footnotetext: https://kaggle.com/datasets/6482c94477918953cec99b03f1330d3afc679e8be2b35113b2463c5688389c18
@@ 1 Introduction
AI models can now classify images, generate text, and create content with high accuracy and fidelity. Consequently, they are used in numerous industries, from national security [62, 5] to healthcare [33, 56]. Given our increasing dependence on AI models in these critical areas, it is critical to interrogate properties of these models and understand their capabilities and limits.
@@
One interesting, well-studied phenomenon is representational similarity across AI models. A large body of work has considered ways in which AI model representations (e.g. their mathematical depictions of real-world inputs) align. Although much of this work has focused on neural networks, recent work has applied similar methods to generative models like large language models (LLMs) and found some evidence of “feature universality” across these models [22].
Although representational alignment is well-documented and novel methods for measuring it are developed frequently, we still do not understand why it happens. Only a few prior works have addressed this question empirically, proposing loss function similarity [45] or inductive bias as possible causal factors [59]. Recent work [22] postulated that task alignment among large-scale AI models will increase their similarity, since as models grow larger they will perform more tasks, and this set of tasks will eventually converge. While compelling, this argument lacks empirical evidence.
@@
We believe another, unexplored factor could also drive model alignment: dataset overlap. Many of today’s massive generative AI models, such as Meta’s LLama, Open AI’s ChatGPT, and Google’s Gemini, are trained on huge swaths of the internet. This design choice follows well-established scaling laws in the AI community, which state that training on more data produces a better model. While massive and relatively cheap to acquire, internet-sourced model training datasets may overlap significantly. For example, many models are trained on subsets of Common Crawl, a massive dataset composed of most of the internet. Many are also trained on Reddit, Wikipedia, and other “known” sources of relatively high quality internet data [42, 3]. It seems not-unreasonable that overlap in training datasets could drive downstream model similarity.
@@
Since limited work explores causes of representational similarity, these two competing hypotheses for model alignment remain unexplored. The task alignment hypothesis of [22] is only ever argued-for theoretically, so no experiments exist. Similarly, the dataset overlap hypothesis not been explored, since the AI research community focuses more on issues of privacy and copyright in large-scale internet data use, not downstream effects on model behaviors.
@@
Our contribution. We propose empirical methods to explore of causes of representational similarity. Building on hypotheses of prior work [22] and known realities of model training, we focus specifically on measuring the effects of dataset overlap and task alignment on representation similarity. To do this, we propose two novel methods, dataset splitting and task splitting, for generating training datasets that control overlap at both the data point and task level, allowing us to quantify how much overlap on these properties affects representation similarity. Along the way, we create a new dataset, called ColorShapeDigit800K, to train models with nuanced task differences, which will be open-sourced upon publication of our paper. We run dataset and task splitting experiments on models ranging from classifiers to small language models and find that:
@@
  • Both dataset and task overlap are positively correlated with increased representational similarity in models. We observe this trend in nearly all models and datasets we evaluate, from image classifiers to VAEs to text generation models.
@@
The rest of the paper proceeds as follows. §2 situates our study in the broader landscape of work on representational similarity. §3 motivates the two causal factors we explore—dataset and task overlap. §4 provides details on our experimental setup, while §5 presents our key findings. §6 discusses limitations and broader impacts of our work.
@@ 2 Related Work
Observations of model similarity. Numerous prior works have demonstrated that machine learning models, even those trained for different tasks, can exhibit similar properties [34, 44, 50, 29, 31, 21, 57], either in their outputs (functional similarity) or their internal behaviors (representational similarity, the focus of this work). Prior work on representational similarity has considered similarities in neuron activation patterns [14, 17] and feature space representations of inputs [31, 21, 44], and numerous metrics have been proposed to measure similarities between models [29, 28, 22, 36, 16, 18, 2, 37, 21]. Extensions of this phenomenon, including alignment between AI model and human representations of the world, is an ongoing area of research (see [57] for a detailed overview).
Causes of model similarity. Prior work has explored factors affecting similarity of computer vision models, such as increases in model and dataset size [29, 52, 22], overall training objectives [8, 32], multi-modality task generalization [22], and model initialization [34]. The main findings show that alignment scales with model and dataset size and that differences in training objective and tasks cause divergent model representations [57]. However, these factors have only been explored at a high level, and little attention has been devoted to untangling direct causes of representational similarity in models. This motivates our work, which empirically studies how two possible causal factors—dataset overlap and task overlap—affect representational similarity.
Downstream effects of representational similarity. Beyond curiosity about causes of model similarity, our work is motivated by concerns about negative effects of unexpected model similarity. Sometimes, similarity is desired and explicitly constructed—e.g. via techniques like model distillation and transfer learning that explicitly pass knowledge from one model to another to expedite learning [48, 6, 57, 54, 11]. However, representation alignment between models may not always be desirable. For example, Wenger and Kenett [63] showed that large language models exhibit homogeneous behaviors on creative tasks. If independently trained models converge to similar representations, their resulting behavior could limit diversity in downstream applications and cause pervasive bias or consistent errors across models. There are numerous applications also in which model diversity matters, such as ensemble learning [38, 64] or multi-agent systems [19]. We believe that studying causes of representational similarity across models could reveal mitigation strategies to address unwanted model convergence.
@@ 3 Setup and Intuition
where $\mathcal{L}$ is the training objective. The model consists of $f$ and $h$ parameterized by $\hat{\theta}_{i}$, where $f$ is the backbone that generates a rich latent-space representation of input $z$ and $h$ is a small head module that transforms the latent vector into human-usable outputs, perhaps classification decisions or generated text. In this work, we are interested in the output of $f$, where $f_{\hat{\theta}_{i}}:\mathbb{D}\rightarrow\mathbb{R}^{n}$ maps an input $z\in\mathbb{D}$ to an $n$-dimensional latent representation of input $z$. Given an input $z\in\mathbb{D}$ and model backbones $f_{\hat{\theta}_{1}}$ and $f_{\hat{\theta}_{2}}$, we can measure the similarity in the latent representations of $f_{\hat{\theta}_{1}}(z)$ and $f_{\hat{\theta}_{2}}(z)$. Figure 1 provides an overview of this end-to-end framework, which forms the backbone of our methodology.
@@
Numerous prior works have measured similarities between $f_{\hat{\theta}_{1}}$ and $f_{\hat{\theta}_{2}}$. However, little work has explored causes of such similarities. In this work, we propose to study the relative effects of two potential causes on downstream representational similarity: dataset overlap and task overlap. Given the wide open nature of this line of inquiry, we believe that focusing on two potential causes is sufficient. Future work should rigorously examine additional causal factors.
@@ 3.1 Dataset overlap
Our exploration of this factor is motivated by the reality that today’s large-scale AI models are trained on overlapping datasets. Due to ever-increasing demands for training data [23], model trainers often turn to internet scrapes as a rich and cheap source of data. From whitepapers, one can easily discern that the scraped training dataset of large models often overlap. For example, GPT [4], Jamba [58], Llama [60], PaLM [7], and Phi [1] are all trained on subsets of CommonCrawl [10], while GPT, Llama, and PaLM are all trained on Wikipedia and Books datasets. This reality motivates our exploration of how dataset overlap drives downstream representational similarity.
Intuitively, a large amount of overlapping content in training datasets $D_{1}$ and $D_{2}$ (both drawn from the same underlying data distribution $\mathbb{D}$) should lead to high similarity in latent representations produced by the two models. Suppose two datasets overlap significantly, i.e., $D_{1}\bigcap D_{2}\approx D_{1}\bigcup D_{2}$, then solving Equation 1 for $D_{1}$ and $D_{2}$ produces models optimized toward the similar underlying input-label patterns in the two datasets. As a result, the backbones should output similar latent representations for the same data. Prior work shows that models trained on disjoint datasets drawn from the same true distribution (e.g. face images) also exhibit significant similarity in latent space representations [55], increasing our confidence in this hypothesis.
@@ 3.2 Task overlap
Prior work speculates that overlap in model tasks, e.g. specific desired model behaviors, will increase downstream model similarity ([22], Section 3.1). Intuitively, this makes sense: two models trained to answer science and math questions will likely be more similar than a model trained to answer science and math questions and a model trained for general-purpose text completion. However, the claims of [22] are not explored empirically. This motivates our quantitative exploration of task overlap as a second potential cause of representational similarity.
@@ 4 Methodology
To test the relative effects of dataset and task overlap on downstream representational similarity, we trained numerous models with varying data and task overlap in their training datasets. This section describes the techniques used to induce dataset and task overlap and methods for model training and evaluation. Figure 2 gives an overview of our task and dataset splitting approaches.
4.1 Dataset overlap
@@
We split datasets by dividing up data points in each partition and fixing a certain amount of partition-level data point overlap between models. Practically, this means we consider two datasets $D_{1}$ and $D_{2}$, each containing data points drawn from the same set of partitions (e.g. classes in image classification). Given a fixed overlap proportion $\alpha\in[0,1]$, we modify the data points in each partition such that for every partition $p$, a proportion $\alpha$ of the data points in $D_{1}$ and $D_{2}$ are shared (i.e., identical), and the remaining proportion $1-\alpha$ of data points are unique to each dataset. This ensures that each partition in $D_{1}$ and $D_{2}$ has an exact overlap proportion of $\alpha$.
4.2 Task overlap
Task splitting with varying dataset overlap. Defining task similarity requires varying the training objective rather than the data points seen during training. For our baseline approach, we control task similarity between models trained on otherwise-identical datasets $D_{1}$ and $D_{2}$ by varying the number of overlapping partitions $p$ between the two datasets. Specifically, we split the data into partitions $P=\{p_{0},p_{1},\dots,p_{n}\}$, where each partition corresponds to a class (or, in the case of language datasets, tokens from different corpora). We construct $D_{1}$ and $D_{2}$ by selecting $K$ total partitions for each, with $\alpha K$ partitions shared between them and $(1-\alpha)K$ partitions unique to each.
@@ 4.2 Task overlap
Figure 3: Illustration of dataset pairs with different levels of task overlap. Datasets with varying levels of task overlap (they all share the same images). The task overlap progressively increases from left to right. Example images and the corresponding labels are shown for each dataset at a particular task overlap.
Task splitting with constant dataset overlap. We construct a new dataset, ColorShapeDigit800K, which maintains a fixed set of data points while varying the task definition. This approach allows us to control training objective similarity simply by modifying dataset labels. ColorShapeDigit800K is constructed from two datasets: geometric shape images from [51] and digit images from [24]. We choose these base datasets because both contain large collections of visually simple and distinct shapes and digits on clean backgrounds, which enable compositional task construction. More complex data with noisier backgrounds would make it more difficult to control task-specific features.
Table 1: Composition of dataset pairs in ColorShapeDigit800K. The labels of $D_{1},D_{2}$ and the number of classes in $D_{1},D_{2}$ for each level of task overlap is provided as <info for $D_{1}$>/<info for $D_{2}$>. S, D, and C refer to shape, digit, and color, respectively. Figure 6 combines levels 2A, 2B, and 2C.
Overlap level 1 2A 2B 2C 3
@@
For our experiments with this task overlap approach, we use identical data points in both $D_{1}$ and $D_{2}$ and progressively adjust the labelling scheme to define increasingly specific tasks (e.g., based on shape, digit, or color). This is illustrated in Figure 3. As more attributes are incorporated into the task definition, the training objectives across $D_{1}$ and $D_{2}$ become more aligned, although the data points never change. This progression of the labelling scheme to consider more attributes is summarized in Table 1, which shows how the complexity in the number of classes increases as more attributes are considered. At the extreme, when all data attributes are used in labelling, the models are trained on identical tasks. This approach allows us to examine how increasing task similarity affects model representations but avoids the confounding variable of increasing dataset overlap. The downside of this approach is that it leverages a different dataset than that used in our dataset overlap and baseline task overlap experiments, potentially rendering results less comparable.

### v2

# Causes and Consequences of Representational Similarity in Machine Learning Models

@@ Abstract
Numerous works have noted similarities in how machine learning models represent the world, even across modalities. Although much effort has been devoted to uncovering properties and metrics on which these models align, surprisingly little work has explored causes of this similarity. To advance this line of inquiry, this work explores how two factors—dataset overlap and task overlap—influence downstream model similarity. We evaluate the effects of both factors through experiments across model sizes and modalities, from small classifiers to large language models. We find that both task and dataset overlap cause higher representational similarity and that combining them provides the strongest effect. Finally, we consider downstream consequences of representational similarity, demonstrating how greater similarity increases vulnerability to transferable adversarial and jailbreak attacks.
@@ 1 Introduction
A large body of work has considered ways in which AI model representations—models’ mathematical depictions of real-world inputs—align (e.g.  Sucholutsky et al. (2024); Kornblith et al. (2019); Klabunde et al. (2023); Raghu et al. (2017); Morcos et al. (2018); Laakso & Cottrell (2000); Nguyen et al. (2021); Dravid et al. (2023); Zou et al. (2023), among many others). Although much of this work has focused on neural networks, recent work found evidence of “feature universality” (a.k.a. representational alignment) across generative models like large language models (LLMs) (Huh et al., 2024). Although representational alignment across AI models is well-documented and novel methods for measuring it are developed frequently, only limited work works has investigated its root causes. Understanding the causes and consequences of this phenomenon may enable more principled interactions with the AI models now pervasive in critical arenas like healthcare and national security (Wang et al., 2022; Caballero & Jenkins, 2025; Ma et al., 2024; Steinberg et al., 2021).
@@
Huh et al. (2024) postulated that task alignment among large-scale AI models will increase their similarity, since as models grow larger they will perform more tasks, and this set of tasks will eventually converge. While compelling, this argument lacks empirical evidence. Another, yet-unexplored factor could also drive model alignment: dataset overlap. Many of today’s generative AI models are trained on huge swaths of the internet, since AI “scaling laws” suggest that training on more data produces a better model (Kaplan et al., 2020). While cheap to acquire, internet-sourced model training datasets may overlap significantly. For example, many models are trained on subsets of Common Crawl, Reddit, and Wikipedia (Naveed et al., 2025; Bommasani et al., 2022). This overlap could drive downstream model similarity.
@@
Since limited work explores causes of representational similarity, these two competing hypotheses for model alignment remain unexplored. The task alignment hypothesis of Huh et al. (2024) is only ever argued-for theoretically, so no experiments exist. Similarly, the dataset overlap hypothesis not been explored, since the AI research community focuses more on issues of privacy and copyright in large-scale internet data use, not downstream effects on model behaviors.
@@
Our contribution. We propose empirical methods to explore of causes of representational similarity. Building on hypotheses of Huh et al. (2024) and known realities of model training, we explore the relative effects of dataset overlap and task overlap on downstream representation similarity. To do this, we propose two novel methods, dataset splitting and task splitting, that generate datasets with controlled overlap at both the data point and task level, allowing us to quantify how overlap on these properties affects representation similarity. Along the way, we create a new dataset, called ColorShapeDigit800K, to train models with nuanced task differences. We run dataset and task splitting experiments on models ranging from classifiers to small language models and find that:
@@
  • Both dataset and task overlap are positively correlated with increased representational similarity in models. We observe this trend in nearly all models and datasets we evaluate, from image classifiers to diffusion to text generation models.
@@
  • Higher dataset/task overlap cause higher vulnerability to transfer adversarial and jailbreaking attacks, indicating a possible downstream consequence of unexamined model similarity.
The rest of the paper proceeds as follows. §2 situates our study in the broader landscape of representational similarity. §3 motivates the two causal factors we explore—dataset and task overlap. §4 provides details on our experimental setup. §5 and §6 present key findings. §7 discusses limitations and broader impacts of our work. Please see code [^2] https://anonymous.4open.science/r/ReprSimCauses-5D26 and dataset ColorShapeDigit800K [^3] https://drive.google.com/file/d/1N1xWKNWm3rD-a7EOmMwQn86pyus8Xn0P/view?usp=share_link.
@@ 2 Related Work
Observations of model similarity. Numerous prior works have demonstrated that AI models, even those trained for different tasks, can exhibit similar properties, either in their outputs (functional similarity) or their internal behaviors (representational similarity, the focus of this work) (Mehrer et al., 2020; Nguyen et al., 2021; Räuker et al., 2023; Kornblith et al., 2019; Li et al., 2015; Hermann & Lampinen, 2020; Sucholutsky et al., 2024). Prior work on representational similarity has considered similarities in neuron activation patterns and feature space representations of inputs (Li et al., 2015; Hermann & Lampinen, 2020; Nguyen et al., 2021; Dravid et al., 2023; Gurnee et al., 2024), and numerous metrics have been proposed to measure similarities between models (Kornblith et al., 2019; Klabunde et al., 2025; Huh et al., 2024; Morcos et al., 2018; Geirhos et al., 2020; Hacohen et al., 2020; Bansal et al., 2021; Moschella et al., 2023; Hermann & Lampinen, 2020). Extensions of this phenomenon, including alignment between AI model and human representations of the world, is an ongoing area of research (see Sucholutsky et al. (2024) for a detailed overview).
Causes of model similarity. Prior work has explored factors affecting similarity of computer vision models, such as increases in model and dataset size (Kornblith et al., 2019; Roeder et al., 2021; Huh et al., 2024), overall training objectives (Ciernik et al., 2025; Lindsay et al., 2022), multi-modal task generalization (Huh et al., 2024), and model initialization (Mehrer et al., 2020). The main findings show that representational alignment scales with model and dataset size and that differences in training objectives and tasks cause divergent model representations (Sucholutsky et al., 2024). However, these factors have only been explored at a high level, and few efforts have been made to untangle direct causes of representational similarity in models. This motivates our work, which empirically studies how two possible causal factors—dataset overlap and task overlap—affect representational similarity.
Downstream effects of model similarity. Beyond curiosity about causes of model similarity, our work is motivated by concerns about negative effects of unexpected model similarity. Sometimes, similarity is desired and explicitly constructed—e.g. via techniques like model distillation and transfer learning that explicitly pass knowledge from one model to another to expedite learning (Phuong & Lampert, 2019; Cho & Hariharan, 2019; Sucholutsky et al., 2024; Roth et al., 2024; DeepSeek-AI et al., 2025). However, representation alignment between models may not always be desirable. A significant body of work has highlighted the risk of “transferable” adversarial attacks, in which a malicious input designed to mislead one model can also mislead a different model with similar feature representations (Zou et al., 2023; Shan et al., 2020; Demontis et al., 2019; Carlini & Wagner, 2017). Finally, Wenger & Kenett (2025) showed that large language models exhibit homogeneous behaviors on creative tasks. If independently trained models converge to similar representations, this could cause pervasive bias or consistent errors across models. This motivates our §6 investigation of possible consequences from representational similarity.
@@ 3 Setup and Intuition
where $\mathcal{L}$ is the training objective. The model consists of $f$ and $h$ parameterized by $\hat{\theta}_{i}$, where $f$ is the backbone that generates a rich latent-space representation of input $z$ and $h$ is a small head module that transforms the latent vector into human-usable outputs, perhaps classification decisions or generated text. This work considers the output of $f$, where $f_{\hat{\theta}_{i}}:\mathbb{D}\rightarrow\mathbb{R}^{n}$ maps an input $z\in\mathbb{D}$ to an $n$-dimensional latent representation of input $z$. Given an input $z\in\mathbb{D}$ and model backbones $f_{\hat{\theta}_{1}}$ and $f_{\hat{\theta}_{2}}$, we can measure the similarity in the latent representations of $f_{\hat{\theta}_{1}}(z)$ and $f_{\hat{\theta}_{2}}(z)$. Figure 1 provides an overview of this end-to-end framework, which forms the backbone of our methodology.
@@ 3.1 Dataset overlap
Our exploration of this factor is motivated by the reality that today’s large-scale AI models are trained on overlapping datasets. Due to ever-increasing demands for training data (Kaplan et al., 2020), model trainers often turn to internet scrapes as a rich and cheap source of data. From whitepapers, one can easily discern that the scraped training dataset of large models often overlap. For example, GPT (Brown et al., 2020), Jamba (Team et al., 2024), Llama (Touvron et al., 2023), PaLM (Chowdhery et al., 2023), and Phi (Abdin et al., 2023) are all trained on subsets of CommonCrawl (Crawl, 2025), while GPT, Llama, and PaLM are all trained on Wikipedia and Books datasets. This reality motivates our exploration of how dataset overlap drives downstream representational similarity.
Intuitively, a large amount of overlapping content in training datasets $D_{1}$ and $D_{2}$ (drawn from the same underlying data distribution $\mathbb{D}$) should lead to high similarity in models’ latent representations. Suppose two datasets overlap significantly, i.e., $D_{1}\bigcap D_{2}\approx D_{1}\bigcup D_{2}$, then solving Equation˜1 for $D_{1}$ and $D_{2}$ produces models optimized toward the similar input-label patterns in the two datasets. As a result, the backbones should output similar latent representations for the same data. Prior work shows that models trained on disjoint datasets drawn from the same true distribution (e.g. faces) also have similar latent space representations (e.g.  (Shan et al., 2020)), bolstering this hypothesis.
@@ 3.2 Task overlap
Prior work speculates that overlap in model tasks, e.g. specific desired model behaviors, will increase downstream model similarity (Huh et al. (2024), Section 3.1). Intuitively, this makes sense: two models trained to answer science and math questions will likely be more similar than a model trained to answer science and math questions and a model trained for general-purpose text completion. However, the claims of Huh et al. (2024) are not explored empirically. This motivates our quantitative exploration of task overlap as a second potential cause of representational similarity.
@@ 4 Methodology
To test the relative effects of dataset/task overlap on representational similarity, we trained numerous pairs of models with varying data and task overlap in their training datasets, following Figure 1. This section describes our techniques to induce dataset and task overlap and methods for model training and evaluation. Figure 2 gives an overview of our task and dataset splitting approaches.
4.1 Dataset and Task overlap
@@
To control task/dataset overlap, we first subdivide training datasets of interest $D_{1},D_{2}$ into partitions. Partitions are formed contextually based on the dataset domain. For example, in image classification, datasets are partitioned by classes, while in text generation, datasets are partitioned by text style (e.g. code generation vs. science question answering vs. conversational text completion).
Dataset overlap. To control dataset overlap, we fix a certain amount of partition-level data point overlap between models. Practically, this means we consider two datasets $D_{1}$ and $D_{2}$, each containing data points drawn from the same set of partitions (e.g. classes in image classification). Given a fixed overlap proportion $\alpha\in[0,1]$, we modify the data points in each partition such that for every partition $p$, a proportion $\alpha$ of the data points in $D_{1}$ and $D_{2}$ are shared (i.e., identical), and the remaining proportion $1-\alpha$ of data points are unique to each dataset.
Task overlap with varying dataset overlap. Controlling task overlap requires varying models’ training objective rather than the data points seen during training. As a baseline, we control task similarity between models trained on otherwise-identical datasets $D_{1}$ and $D_{2}$ by varying the number of overlapping partitions $p$ between the two datasets. Specifically, we split the data into partitions $P=\{p_{0},p_{1},\dots,p_{n}\}$, where each partition corresponds to a class (or, in the case of language datasets, tokens from different corpora). We construct $D_{1}$ and $D_{2}$ by selecting $K$ total partitions for each, with $\alpha K$ partitions shared between them and $(1-\alpha)K$ partitions unique to each.
@@
Figure 3: Illustration of task partitioning. Datasets with varying levels of task overlap (they all share the same images). The task overlap progressively increases from left to right. Example images and the corresponding labels are shown for each dataset at a particular task overlap.
Task splitting with constant dataset overlap. We construct a new dataset, ColorShapeDigit800K, which maintains a fixed set of data points while varying the task definition. This approach allows us to control training objective similarity simply by modifying dataset labels. ColorShapeDigit800K is constructed from two datasets: geometric shape images from rivaldo (2022) and digit images from Kapoor (2021). We choose these base datasets because both contain large collections of visually simple and distinct shapes and digits on clean backgrounds, which enable compositional task construction.
Table 1: Composition of task partitions in ColorShapeDigit800K. The labels of $D_{1},D_{2}$ and the number of classes in $D_{1},D_{2}$ for each partition is provided as <info for $D_{1}$>/<info for $D_{2}$>. S, D, and C refer to shape, digit, and color, respectively. Figure˜6 combines partitions 2A, 2B, and 2C.
Partition 1 2A 2B 2C 3
@@
For our experiments with this task overlap approach, we use identical data points in both $D_{1}$ and $D_{2}$ and progressively adjust the labelling scheme to define increasingly specific tasks (e.g., based on shape, digit, or color). This is illustrated in Figure 3. As more attributes are incorporated into the task definition, the training objectives across $D_{1}$ and $D_{2}$ become more aligned, although the data points never change. This progression of the labelling scheme to consider more attributes is summarized in Table 1, which shows how the complexity in the number of classes increases as more attributes are considered. When all data attributes are used in labelling, the models are trained on identical tasks. This approach allows us to examine how increasing task similarity without the confounding variable of increasing dataset overlap affects model representations.